OpenStudy (anonymous):

The sum of the digits of a certain two-digit number is 7. Reversing its digits increases the number by 9. What is the number?

OpenStudy (anonymous):

There are actually two ways to solve this problem.

OpenStudy (anonymous):

x+y=7

OpenStudy (anonymous):

The first way is to derive two additional equations. If x is the 10s digit and y is the ones digit then 10x plus y equals the number.

OpenStudy (anonymous):

10x+y=10

OpenStudy (anonymous):

Now, if we reverse the digits, y becomes the 10s digit and x becomes the ones digit and the number is increased by 9, so: x+10y=n+9

OpenStudy (anonymous):

Add -9 to both sides of this last equation to get x+10y-p=n . Now we have two things that equal n so we can set these two expressions equal to each other

OpenStudy (anonymous):

x+10y-9=n**

OpenStudy (anonymous):

10x+y=x+10y-9

OpenStudy (anonymous):

9x-9y=-9

OpenStudy (anonymous):

x-y=1

OpenStudy (anonymous):

Add this last equation to your very first equation (x+y=7) term by term:

OpenStudy (anonymous):

2x+0y=6

OpenStudy (anonymous):

x=3 from 3+y=7 we get y=4 therefore the number is 34

OpenStudy (anonymous):

oh ok thats what i got thanks!

OpenStudy (anonymous):

np